OpenStudy (anonymous):

PLEASEEEE Which of the following systems of equations has infinitely many solutions? 3x - 3y = 11; 6y = 9x - 22 5y = -9x - 6; -27x + 15y = 15 4x - 7y = 7; -14y = 8x - 14 -4x + 2 = 3y; 12y - 8 = -16x

OpenStudy (allank):

The last option because the first and second equations are the same. To be able to see that, rearrange the second one like this: -16x+8=12y then divide through by 4 to get : -4x+2=3y

OpenStudy (anonymous):

good question! so let me try if the determinant of a system is zero then the system have infinite many solution let me solve first one i hope you can do the rest first of all the equations must be standard form ax+by=c form we have 3x - 3y = 11; .............1 6y = 9x - 22 ................2 write the second equation in standard form 9x-6y=22 so we have two equations in standard form 3x - 3y = 11; 9x-6y = 22; write in matrix form \[\Large A=\left[\begin{matrix}3 & -3 \\ 9 & -6\end{matrix}\right]\] now taking determinant \[\Large \det(A)=(3)(-6)-(9)(-3)=9\] since \[\Large \det(A)\neq0\] so the system does not have infinite many solutions! can you check the rest in the same way???

OpenStudy (anonymous):

so its the last one?

OpenStudy (anonymous):

@mc32141 check the determinant of the last one if it is zero then yes it is !

OpenStudy (allank):

@sami-21 Hadn't thought about it that way. nice. :) And yes, @mc32141 you should get the determinant as zero.

OpenStudy (anonymous):

@allank i think it is easy way :) because if the two rows in determinant will gets similar then it will be zero.

OpenStudy (anonymous):

@mc32141 did you get it?